# Text2PDE: Latent Diffusion Models for Accessible Physics Simulation

### v1

@@ 1 Introduction
Neural PDE solvers are an exciting new class of physics solvers that have the potential to improve many aspects of conventional numerical solvers. Initial works have proposed various architectures to accomplish this, such as graph-based (Li & Farimani, 2022; Battaglia et al., 2016), physics-informed (Raissi et al., 2019), or convolutional approaches (Thuerey et al., 2020). Subsequent work on developing neural operators (Li et al., 2021; Lu et al., 2021; Kovachki et al., 2023) established them as powerful physics approximators that can quickly and accurately predict PDEs, and recent work has focused on improving many of their different aspects. These advances have established different models that can adapt to irregular grids (Brandstetter et al., 2023; Li et al., 2023a; c; Wu et al., 2024), speed up model training (Li et al., 2024a; Alkin et al., 2024; Tran et al., 2023; Li et al., 2024b; 2023b), achieve more accurate solutions (Lippe et al., 2023; Li et al., 2023d), or generalize to a wide variety of parameters (McCabe et al., 2023; Hao et al., 2024; Herde et al., 2024; Zhou et al., 2024).
Figure 1: We introduce latent diffusion models for physics simulation, with the remarkable ability of generating an entire PDE rollout from a text prompt. Two generated solutions are displayed with their model inputs.
Despite these advances, there are still many factors that limit the practical adoption of neural PDE surrogates. This is a challenging task for any emerging field: practitioners must identify novel research, train a model for their own needs, and be able to accurately deploy models beyond established benchmarks. As a field, amazing progress is being made to address this through advancing accuracy and generalizability, but this alone may not be enough. In particular, the study of neural PDE surrogates is highly interdisciplinary and fragmented. Practitioners rely on physics and numerical analysis expertise to solve PDEs, while novel research in neural solvers leverages highly specialized deep learning knowledge and practices. This makes it challenging to train and deploy neural PDE surrogates, as engineers need to learn new domain expertise to train or query novel architectures for their needs.
@@
In this work, we take a step towards addressing this by creating more accessible physics simulators that are still accurate and efficient to train, which ensures suitable performance and feasible deployment. We accomplish this by introducing language as a novel modality for interacting with physics solvers, creating text2PDE models. Language is extremely general: while different physical phenomena usually require problem-specific meshing and solvers, humans constantly observe and interact with physics and, as a result, build a common understanding and vocabulary for describing these phenomena. This lowers the barrier to entry in adopting neural surrogates: rather than needing deep learning expertise to properly query neural surrogates, practitioners can easily express ideas in language to prompt a physics model. Indeed, the popularity of text2image models is partly due to the ease of expressing visual ideas in language, rather than, for example, needing the computer vision expertise to query a class-conditional image generation model with an ImageNet label.
Beyond improving accessibility and interpretability, language can also have technical benefits. As long as the physical phenomena can be described in text, its domain specialties can be abstracted away, such as the discretization, geometries, boundaries, or physical variables. Rather than needing separate methods based on different physics expertise, text2PDE models only need to work with a single modality of language. In addition, language can be very compact; rather than needing raw physical data to specify a simulation, this information can be distilled into rich, dense textual descriptions. However, this presents unique challenges. Since text can abstractly describe a broad set of physics, the text2PDE model must generate arbitrarily discretized results. Furthermore, language can be imprecise in describing many physics problems. Lastly, producing high-dimensional spatio-temporal physics simulations from condensed text information is challenging.
@@
To address these challenges, we introduce methods to use latent diffusion models to generate PDE solutions. Within this framework, we generate an entire solution trajectory at once to avoid autoregressive error propagation and improve accuracy. Furthermore, we develop a novel autoencoder strategy to efficiently train and query models as well as to adapt latent diffusion to arbitrary discretizations. Moreover, we investigate conditioning on either text or physics modalities, such as the initial frame of a solution, to improve the precision of generated results. Lastly, we scale latent diffusion models up to nearly 3 billion parameters to show their ability to accurately generate complex outputs from compact inputs. We believe that this scalability, accuracy, and usability can unlock generative models that can bring neural PDE surrogates closer to practical adoption. To further this objective, all code, datasets, and pretrained models are released at: https://github.com/anthonyzhou-1/ldm_pdes
@@ 2.1 Problem Setup
Numerical Solvers
Conventional numerical solvers approach solving time-dependent PDEs by spatially discretizing the solution $\mathbf{u}$ at a timestep $t$ to obtain $\mathbf{u}(t,\mathbf{x_{m}})$ collocated on $M$ mesh points $\mathbf{x}_{m}\subset\Omega$ for $m\in\{0,1,\ldots,M\}$. To advance the solution in time, the spatial derivatives ($\partial_{x}\mathbf{u},\partial_{xx}\mathbf{u},\ldots$) are approximated to compute an estimate of $\partial_{t}\mathbf{u}(t,\mathbf{x}_{m})$ and used to advance the current solution in time: $\mathbf{u}(t,\mathbf{x_{m}})\rightarrow\mathbf{u}(t+\Delta t,\mathbf{x_{m}})$, also known as the method of lines (Schiesser, 2012). Many schemes have been developed to approximate spatial derivatives or advance solutions in time, which have varying complexities and accuracies (Butcher, 1987).
Neural Solvers
Neural solvers aim to solve time-dependent PDEs by using the solution at a current timestep to predict the solution at a future timestep $\mathbf{u}(t+\Delta t,\mathbf{x_{m}})=f_{\theta}(\mathbf{u}(t,\mathbf{x_{m}}))$. Many variants of this setup exist, whether it is to use multiple input timesteps or to predict multiple output timesteps. Despite these variants, the dominant approach still follows numerical solvers by autoregressively advancing the solution in time, with certain notable exceptions (Lienen et al., 2024; Du et al., 2024).
Autoregressive solvers are a natural choice for neural surrogates; they follow numerical solver precedents, as well as reflect the Markovian nature of time-dependent PDEs. However, this paradigm has many potential drawbacks, namely the stability of autoregressive solvers (Lippe et al., 2023). Although many approaches have been developed to mitigate autoregressive error accumulation (Brandstetter et al., 2023), the fundamental problem still exists: complex PDEs require small timescales to accurately resolve physical phenomena, yet this requires longer rollouts which increase error accumulation (Lozano-Durán & Jiménez, 2014; Lienen et al., 2024). To address this issue, we leverage spatio-temporal diffusion to directly predict an entire solution trajectory. This becomes a generative, rather than autoregressive, problem conditioned on information about the desired PDE solution.
@@ 2.2 Spatio-temporal Diffusion for PDEs
In the context of PDEs, this model would non-deterministically sample a full rollout $\mathbf{u}$ given its initial and boundary conditions. Since $\mathbf{u}$ can be extremely high-dimensional, we perform the forward and reverse diffusion process in a learned latent space. More specifically, given a solution $\mathbf{u}\in\mathbb{R}^{T\times M\times d_{p}}$, where $T$ represents the temporal resolution, $M$ represents the spatial resolution, and $d_{p}$ is the physical dimension, an encoder $\mathcal{E}$ can embed $\mathbf{u}$ into a latent vector $\mathbf{z}=\mathcal{E}(\mathbf{u})$. Importantly, $\mathbf{z}$ is in a compressed latent space $\mathbb{R}^{L\times d_{l}}$, where $L$ is a latent spatio-temporal resolution smaller than $T\times M$, and $d_{l}$ is the latent dimension. To project latent vectors back to the physical space, a decoder $\mathcal{D}$ is used to reconstruct a sample $\mathbf{u}=\mathcal{D}(\mathbf{z})$.
@@ 3.1 Autoencoders for PDE Data
PDE data can vary widely between equations and use cases; solutions can be discretized on a uniform or irregular mesh or contain varying physical quantities (e.g., pressure, velocity, density, etc.). In particular, practical physics applications often discretize solutions on an unstructured mesh, which aids in improving accuracy at regions of interest. This can be difficult for autoencoders to accommodate, and indeed much of the previous work is focused on compressing uniformly structured images, videos, or PDE solutions (He et al., 2021; Feichtenhofer et al., 2022; Zhou & Farimani, 2024).
@@
Following the problem setup, at every arbitrary timestep $t\in[0,T]$ we can consider $M$ arbitrary coordinates $\mathbf{x}_{m}\subset\Omega$ for $m\in\{0,1,\ldots,M\}$, where a solution vector $\mathbf{u}(t,\mathbf{x}_{m})\in\mathbb{R}^{d_{p}}$ is also defined. This framework captures solutions defined at any point in space and time. We map input timesteps $t$, coordinates $\mathbf{x}_{m}$, and solutions $\mathbf{u}(t,\mathbf{x}_{m})$ onto $T_{l}$ uniform timesteps $t_{l}\in[0,T_{l}]$, $M_{l}$ uniform grid coordinates $\mathbf{x}_{l}\subset\Omega$ for $l\in\{0,1,\ldots,M_{l}\}$, and latent vectors $\mathbf{q}(t_{l},\mathbf{x}_{l})\in\mathbb{R}^{d_{p}}$ at each uniform spatio-temporal coordinate. This is done with the following kernel integral:
@@
where $\mu(\mathbf{y}_{b})$ is the Riemann sum weight for every point $\mathbf{y}_{b}$ in the ball $B_{r}(\mathbf{y}_{l})$. To define and evaluate a latent grid, we partition the time and spatial axes into $T_{l}$ and $M_{l}$ equispaced points, resulting in $T_{l}\times M_{l}$ coordinates $\mathbf{y}_{l}$, each of which needs to have the kernel integral evaluated for it. This kernel integral essentially aggregates neighboring physical solutions for every latent coordinate.
@@
Using this method allows our autoencoder to benefit from architecture unification, where best practices from separate domains can benefit PDE applications. In particular, we can leverage variational or vector quantized latent spaces (Kingma & Welling, 2022; Esser et al., 2021) to avoid arbitrarily high-variance latent spaces during training. Furthermore, to improve reconstruction quality, the autoencoder can benefit from incorporating a generative adversarial network (GAN) or perceptual loss metric (Rombach et al., 2022; Zhang et al., 2018). We leave further details and ablation studies of these modifications to Appendix B.1. As a summary, we find that GANs and perceptual guidance can in certain cases improve reconstruction performance, but for simplicity, we omit them in our main results. Furthermore, we find that highly regularized latent spaces reduce downstream diffusion performance due to lower reconstruction accuracy, and that diffusion backbones are sufficiently parameterized to model higher-variance latent spaces. As a result, our main results use a small Kullback–Leibler (KL) penalty during autoencoder training.
@@
To move from the latent grid to the physical mesh, the kernel integral can be reversed. Specifically, given a decoded latent grid $\mathbf{q}_{d}(\mathbf{y}_{l})$, the decoded solution can be calculated by the Riemann sum: $\mathbf{u}_{d}(\mathbf{y})=\sum_{b}^{M_{b}}\kappa(\mathbf{y},\mathbf{y}_{b})\mathbf{q}_{d}(\mathbf{y}_{b})\mu(\mathbf{y}_{b})$. Note that in this case, the sum is evaluated over the ball $B_{r}(\mathbf{y})$, and the coordinates $\mathbf{y}_{b}$ are defined as the latent grid points $\mathbf{y}_{l}$ that lie within $B_{r}(\mathbf{y})$. This reverse process essentially aggregates neighboring latent vectors for every physical coordinate. Lastly, for PDE problems that are uniformly discretized or resampled onto a uniform grid, the kernel integration can be omitted and a CNN autoencoder can be directly used.
@@ 3.2 Latent Diffusion
Once a latent space has been learned, we consider freezing the encoder and decoder to learn the latent distribution $p(\mathbf{z})$. Specifically, we train a denoising model $\epsilon_{\theta}(\mathbf{z}_{n},n)$, where $n\in\{0,1,\ldots,N\}$ is an intermediate step within the Gaussian noising process of length $N$, and $\mathbf{z}_{n}$ is the noised latent at step $n$  [^1] We deviate from the usual notation of $t$ representing the noising timestep to avoid overloading it with $t$ which is previously defined as physical time of the PDE solution $\mathbf{u}(t,\mathbf{x})$.. This model is trained by minimizing a reweighted variational lower bound on $p(\mathbf{z})$:
$L=\mathbb{E}_{\mathbf{z},\epsilon\sim\mathcal{N}(0,\mathbf{I}),n}\left[\left\lVert\epsilon-\epsilon_{\theta}(\mathbf{z}_{n},n)\right\rVert_{2}^{2}\right]$ (7)
Since the latent $\mathbf{z_{0}}$ is encoded from a uniform latent grid, we can leverage a Unet to parameterize $\epsilon_{\theta}$ (Ronneberger et al., 2015). However, recent work by Peebles & Xie (2023) suggests that this inductive bias is not strictly necessary, and we also evaluate the use of a diffusion transformer (DiT). Empirically, we find that the DiT backbone outperforms the Unet backbone; the main results are reported with a DiT implementation, but a full comparison can be found in Appendix C.2.
Additionally, recent advances in diffusion frameworks have proposed many modifications to the denoising process, the most straightforward of which is scaling the latent space (Rombach et al., 2022). Furthermore, Nichol & Dhariwal (2021) has shown the potential benefits of using a cosine noise scheduler or learning the variance $\Sigma_{\theta}(\mathbf{z}_{n},n)$ of the reverse process. Lastly, Salimans & Ho (2022) introduced reparameterizing the reverse process to predict the velocity $v=\sqrt{\bar{\alpha}_{n}}\epsilon-\sqrt{1-\bar{\alpha}_{n}}\mathbf{x}_{0}$, with $\bar{\alpha}_{n}$ defined as in Ho et al. (2020). We implement and evaluate the effects of these modifications; for interested readers, the findings can be found in Appendix C.1. In general, we find that these modifications can have a large effect on the performance of our model; however, we leave optimizing the diffusion process to future work. For our main results, we scale the latent space, use a linear noise schedule, do not learn the variance, and use an $\epsilon$ parameterization.

### v2

@@ 1 Introduction
Neural PDE solvers are an exciting new class of physics solvers that have the potential to improve many aspects of conventional numerical solvers. Initial works have proposed various architectures to accomplish this, such as graph-based (Li & Farimani, 2022; Battaglia et al., 2016), physics-informed (Raissi et al., 2019), or convolutional approaches (Thuerey et al., 2020). Subsequent work on developing neural operators (Li et al., 2021; Lu et al., 2021; Kovachki et al., 2023) established them as powerful physics approximators that can quickly and accurately predict PDEs, and recent work has focused on improving many of their different aspects. These advances have established different models that can adapt to irregular grids (Brandstetter et al., 2023; Li et al., 2023a; c; Wu et al., 2024), speed up model training (Li et al., 2024a; Alkin et al., 2024; Tran et al., 2023; Li et al., 2024b; 2023b), achieve more accurate solutions (Lippe et al., 2023), or generalize to a wide variety of parameters (McCabe et al., 2023; Hao et al., 2024; Herde et al., 2024; Zhou et al., 2024).
Figure 1: We introduce latent diffusion models for physics simulation, with the remarkable ability of generating an entire PDE rollout from a text prompt. Three generated solutions are displayed with their model inputs.
Despite these advances, there are still many factors that limit the practical adoption of neural PDE solvers. Great progress has been made in addressing the accuracy and generalizability of neural solvers, however, deploying these advances will require further work in developing new engineering tools. Although neural solvers have become highly capable, they have also become highly specialized; increasingly, the practical use of neural solvers is limited by their inaccessibility and the complexity of the field. Deep learning has already enabled vision and language tools designed with productivity and usability in mind; is there a way we can create similar tools for engineers?
@@
In this work, we take a step towards creating more accessible physics simulators by introducing language as a novel modality to interact with physics solvers, creating text2PDE models. This has both benefits and drawbacks. One potential benefit is that language is extremely general: while different physical phenomena usually require problem-specific meshing and solvers, humans constantly observe and interact with physics and, as a result, build a common understanding and vocabulary for describing these phenomena. This could reduce the barrier to entry for adopting neural solvers by allowing practitioners to express ideas in language to prompt a physics model. Indeed, the popularity of text2image models is partly due to the ease of expressing visual ideas in language, rather than, for example, needing the computer vision expertise to query a class-conditional image generation model with an ImageNet label. Beyond improving accessibility, language can also have technical benefits. As long as the physical phenomena can be described in text, its domain specialties can be abstracted away, such as the discretization, geometries, or boundaries; this allows models to work with just a single modality. In addition, language can be compact: rather than needing raw physical data to specify a simulation, this information can be distilled into dense textual descriptions.
However, text2PDE models present unique challenges. Since text can describe a broad set of physics, the text2PDE model must generate arbitrarily discretized results. Furthermore, producing high-dimensional spatio-temporal physics simulations from a low-dimensional text prompt can be challenging. Lastly, language can be imprecise in describing many physics problems, and in certain cases, specifying exact initial conditions may be preferred. Indeed, we envision text2PDE models not as a replacement for current methods, but rather as an additional tool for engineers to leverage. We hope to empower engineers with a diverse toolkit, where different neural solvers and even numerical methods can be used together throughout the design cycle to better ideate, simulate, and realize engineering outcomes.
@@
To work towards this goal, we introduce methods to use latent diffusion models to generate PDE solutions. Within this framework, we generate an entire solution trajectory at once to avoid autoregressive error propagation and improve accuracy. We also develop a novel autoencoder strategy to efficiently train/query models and to adapt latent diffusion to arbitrary discretizations. We also investigate conditioning on either text or physics modalities, such as the initial frame of a solution, to switch between an interpretable interface and a more precise one. Lastly, we scale latent diffusion models up to nearly 3 billion parameters to show their ability to accurately generate complex outputs from compact inputs. We believe that this scalability, accuracy, and usability can unlock generative models that can bring neural PDE solvers closer to practical adoption. To further this objective, all code, datasets, and pretrained models are released at: https://github.com/anthonyzhou-1/ldm_pdes
@@ 2.1 Problem Setup
In general, the solution $\mathbf{u}$ is discretized at a timestep $t$ to obtain $\mathbf{u}(t,\mathbf{x_{m}})$ collocated on $M$ mesh points $\mathbf{x}_{m}\subset\Omega$ for $m\in\{0,1,\ldots,M\}$. Neural solvers aim to solve time-dependent PDEs by using the solution at a current timestep to predict the solution at a future timestep $\mathbf{u}(t+\Delta t,\mathbf{x_{m}})=f_{\theta}(\mathbf{u}(t,\mathbf{x_{m}}))$. Many variants of this setup exist, whether it is to use multiple input timesteps or to predict multiple output timesteps. Despite these variants, the dominant approach still follows numerical solvers by autoregressively advancing the solution in time, with certain notable exceptions (Lienen et al., 2024; Du et al., 2024).
Autoregressive solvers are a natural choice for neural solvers; they follow numerical solver precedents, as well as reflect the Markovian nature of time-dependent PDEs. However, this paradigm has many potential drawbacks, namely the stability of autoregressive solvers (Lippe et al., 2023). Although many approaches have been developed to mitigate autoregressive error accumulation (Brandstetter et al., 2023), the fundamental problem still exists: complex PDEs require small timescales to accurately resolve physical phenomena, yet this requires longer rollouts which increase error accumulation (Lozano-Durán & Jiménez, 2014; Lienen et al., 2024). To address this issue, we leverage spatio-temporal diffusion to directly predict an entire solution trajectory. This becomes a generative, rather than autoregressive, problem conditioned on information about the desired PDE solution.
@@ 2.2 Spatio-temporal Diffusion for PDEs
Since $\mathbf{u}$ can be extremely high-dimensional, we perform the forward and reverse diffusion process in a learned latent space. More specifically, given a solution $\mathbf{u}\in\mathbb{R}^{T\times M\times d_{p}}$, where $T$ represents the temporal resolution, $M$ represents the spatial resolution, and $d_{p}$ is the physical dimension, an encoder $\mathcal{E}$ can embed $\mathbf{u}$ into a latent vector $\mathbf{z}=\mathcal{E}(\mathbf{u})$. Importantly, $\mathbf{z}$ is in a compressed latent space $\mathbb{R}^{L\times d_{l}}$, where $L$ is a latent spatio-temporal resolution smaller than $T\times M$, and $d_{l}$ is the latent dimension. To project latent vectors back to the physical space, a decoder $\mathcal{D}$ is used to reconstruct a sample $\mathbf{u}=\mathcal{D}(\mathbf{z})$.
@@ 3.1 Autoencoders for PDE Data
Practical physics applications often discretize solutions on an unstructured mesh, which helps improve the accuracy at regions of interest. This can be difficult for conventional autoencoders to accommodate, and indeed much of the previous work is focused on compressing uniformly structured images, videos, or PDE solutions (He et al., 2021; Feichtenhofer et al., 2022; Zhou & Farimani, 2024). Therefore, we propose an autoencoder that can compress data given at arbitrary collocation points and reconstruct them at arbitrary query points to extend diffusion models to unstructured data.
@@
Following the problem setup, at every timestep $t\in[0,T]$ we can consider $M$ coordinates $\mathbf{x}_{m}\subset\Omega$ for $m\in\{0,1,\ldots,M\}$, where a solution vector $\mathbf{u}(t,\mathbf{x}_{m})\in\mathbb{R}^{d_{p}}$ is also defined. This framework captures solutions defined at any point in space and time. We map input timesteps $t$, coordinates $\mathbf{x}_{m}$, and solutions $\mathbf{u}(t,\mathbf{x}_{m})$ onto $T_{l}$ uniform timesteps $t_{l}\in[0,T_{l}]$, $M_{l}$ uniform grid coordinates $\mathbf{x}_{l}\subset\Omega$ for $l\in\{0,1,\ldots,M_{l}\}$, and latent vectors $\mathbf{q}(t_{l},\mathbf{x}_{l})\in\mathbb{R}^{d_{p}}$ at each uniform spatio-temporal coordinate. This is done with the following kernel integral:
@@
where $\mu(\mathbf{y}_{b})$ is the Riemann sum weight for every point $\mathbf{y}_{b}$ in the ball $B_{r}(\mathbf{y}_{l})$. This kernel integral essentially aggregates neighboring physical solutions for every latent coordinate and has the nice property of approximating an operator mapping from the input function $\mathbf{u}(\mathbf{y})$ to the latent function $\mathbf{q}(\mathbf{y}_{l})$. Additional details and visualizations of this kernel integral are provided in Appendix B.3.
@@
Using this method allows our autoencoder to benefit from architecture unification, where best practices from separate domains can benefit PDE applications. In particular, we can leverage variational or vector quantized latent spaces (Kingma & Welling, 2022; Esser et al., 2021) to avoid arbitrarily high-variance latent spaces during training. Furthermore, to improve reconstruction quality, the autoencoder can benefit from incorporating a generative adversarial network (GAN) or perceptual loss metric (Rombach et al., 2022; Zhang et al., 2018). We leave further details and ablation studies of these modifications to Appendix B.1. As a summary, we find that GANs and perceptual guidance can in certain cases improve reconstruction performance; however for simplicity and to ensure the broad applicability of our framework, we omit them in our main results since they require extensive optimization and are highly problem-dependent. Furthermore, we find that highly regularized latent spaces reduce downstream diffusion performance due to lower reconstruction accuracy, and that diffusion backbones are sufficiently parameterized to model higher-variance latent spaces. As a result, our main results use a small Kullback–Leibler (KL) penalty during autoencoder training.
@@
To move from the latent grid to the physical mesh, the kernel integral can be reversed. Specifically, given a decoded latent grid $\mathbf{q}_{d}(\mathbf{y}_{l})$, the decoded solution can be calculated by the Riemann sum: $\mathbf{u}_{d}(\mathbf{y})=\sum_{b}^{M_{b}}\kappa(\mathbf{y},\mathbf{y}_{b})\mathbf{q}_{d}(\mathbf{y}_{b})\mu(\mathbf{y}_{b})$. Note that in this case, the sum is evaluated over the ball $B_{r}(\mathbf{y})$, and the coordinates $\mathbf{y}_{b}$ are defined as the latent grid points $\mathbf{y}_{l}$ that lie within $B_{r}(\mathbf{y})$. This reverse process essentially aggregates neighboring latent vectors for every physical coordinate, and also has the property of being discretization-invariant. In particular, the latent solution can be decoded onto an arbitrary mesh by evaluating the Riemann sum at arbitrary query points $\mathbf{y}$, shown in Figure 8. Lastly, for PDE problems that are uniformly discretized, the kernel integration can be omitted and a CNN autoencoder can be directly used.
@@ 3.2 Latent Diffusion
Once a latent space has been learned, we consider freezing the encoder and decoder to learn the latent distribution $p(\mathbf{z})$. Specifically, we train a denoising model $\epsilon_{\theta}(\mathbf{z}_{n},n)$, where $n\in\{0,1,\ldots,N\}$ is an intermediate step within the Gaussian noising process of length $N$, and $\mathbf{z}_{n}$ is the noised latent at step $n$  [^1] We deviate from the usual notation of $t$ representing the noising timestep to avoid overloading it with $t$ which is previously defined as physical time of the PDE solution $\mathbf{u}(t,\mathbf{x})$.. This model is trained by minimizing a reweighted variational lower bound on $p(\mathbf{z})$: $L=\mathbb{E}_{\mathbf{z},\epsilon\sim\mathcal{N}(0,\mathbf{I}),n}\left\lVert\epsilon-\epsilon_{\theta}(\mathbf{z}_{n},n)\right\rVert_{2}^{2}$. Since the latent $\mathbf{z_{0}}$ is encoded from a uniform latent grid, we can leverage a Unet to parameterize $\epsilon_{\theta}$ (Ronneberger et al., 2015). However, recent work by Peebles & Xie (2023) suggests that this inductive bias is not necessary and we evaluate the use of a diffusion transformer (DiT). We find that the DiT backbone outperforms the Unet backbone; the main results are reported with a DiT implementation, but a comparison can be found in Appendix C.2.
Additionally, recent advances in diffusion frameworks have proposed many modifications to the denoising process, the most straightforward of which is scaling the latent space (Rombach et al., 2022). Furthermore, Nichol & Dhariwal (2021) has shown the potential benefits of using a cosine noise scheduler or learning the variance $\Sigma_{\theta}(\mathbf{z}_{n},n)$ of the reverse process. Lastly, Salimans & Ho (2022) introduced reparameterizing the reverse process to predict the velocity $v=\sqrt{\bar{\alpha}_{n}}\epsilon-\sqrt{1-\bar{\alpha}_{n}}\mathbf{x}_{0}$, with $\bar{\alpha}_{n}$ defined as in Ho et al. (2020). We implement and evaluate the effects of these modifications; for interested readers, the findings can be found in Appendix C.1. In general, we find that these modifications can have varying effects on the performance of our model; however, we leave optimizing the diffusion process to future work. For our main results, we scale the latent space, use a linear noise schedule, do not learn the variance, and use an $\epsilon$ parameterization.